OpenStudy (anonymous):

Write the equation of a line in slope-intercept form passing through (4, 5) and (6, 8) A. y equals two-thirds x plus seven-thirds B. y equals three-halves x plus 11 C. y equals three-halves x minus 1 D. y equals three-halves minus 8

OpenStudy (jackellyn):

First we find the slope using the formula: m= y2-y1/x2-x1 where y2=8 y1=5 x2=6 and x1=4 Solve for m. 8-5 3 m=----- = --- 6-4 2 y=mx+b 5=(3/2)(4) + b (Substitute m, x, and y) 5 = 12/2 +b (Solve for b.) 5 = 6 +b 5-6 = b -1 = b y=3/2x -1 (Rewrite the equation.)

OpenStudy (anonymous):

thanks

OpenStudy (jackellyn):

np :)